The length and width of a rectangle are in a ratio of 5 to 4, and the area is 500 square meters. Find the two dimensions.

Found 2 solutions by cleomenius, ewatrrr:
Answer by cleomenius(959)

5x * 4x = 20x^2 = 500
x^2 = 25
x = 5
25 m, 20 m= 500sq m
Cleomenius

Answer by ewatrrr(24785)

 
Hi,
L/W = 5/4 OR L = (5/4)W
A = L*W = (5/4)W*W = 500m^2
W^2 = 400
W = 20 (tossing out negative solution for the # of meters
Width is 20m and length is 25m |(5/4)W
CHECKING our Answer***
20m*25m = 500m^2